OpenStudy (anonymous):

Find the solution to the following problem in correct scientific notation: (3.4 × 10^100) × (2.1 × 10^99) A. 3.61 × 10^100 B. 2.44 × 10^99 C. 7.14 × 10^199 D. 7.14 × 10^1 E. 7.14 × 10^9,900

OpenStudy (anonymous):

Is D the correct answer?

OpenStudy (mayankdevnani):

no!! you are wrong

OpenStudy (mayankdevnani):

please show your work

OpenStudy (mayankdevnani):

@RH

OpenStudy (mayankdevnani):

okkk... first divide: \[\huge \frac{3.4 × 10^{100}}{2.1 × 10^{99}}\] so what you get??

OpenStudy (anonymous):

Sorry its multiplying not dividing, so \[(3.4 x 10^{100}) x (2.1 x 10^{99})\]

OpenStudy (mayankdevnani):

oh!!! yeah!!! sorry.. so first multiply : \[\huge 3.4 × 10^{100} \times 2.1 × 10^{99}\] so what you get?

OpenStudy (anonymous):

7.14 x 10 (I don't know how to multiply with the 10 ?)

OpenStudy (mayankdevnani):

 understood?

thomaster (thomaster):

It's not that difficult :P \(\Large a^m*a^n = a^{m+n}\)

OpenStudy (anonymous):

So C is the correct answer?

thomaster (thomaster):

\(\huge 3.4*2.1=?\)

OpenStudy (anonymous):

7.14

thomaster (thomaster):

So \(\LARGE 3.4^{100}*2.1^{99} = 7.14^{100+99}\)

OpenStudy (mayankdevnani):

always remember that: if the 10th power is multiplied by any 10th power then always be added its power like that: \[\huge 10^5+10^6=10^{5+6}=10^{11}\] and when 10th power is divided by any 10th power then the power will be subtracted like that: \[\huge \frac{10^6}{10^5}=10^{6-5}=10^1\]

OpenStudy (mayankdevnani):

i hope you got it!! @RH

OpenStudy (anonymous):

Yes!!! Thank you very much!

OpenStudy (mayankdevnani):

welcome:)

thomaster (thomaster):

wow i just realize i forgot the 10's :O \(\LARGE 3.4\underline{*10}^{100}\ * ~~ 2.1\underline{*10}^{99} = 7.14\underline{*10}^{100+99}\) My apologies if it made it confusing

OpenStudy (anonymous):

No problem! Thanks a lot :)